OpenStudy (anonymous):

Factor the trinomial below. 4x2 + 23x - 6 Which of the following binomials is a factor of the trinomial?

OpenStudy (anonymous):

4x2 + 24x -x - 6 =4x(x+6) -1(x+6)=(x+6)(4x-1)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

welcome dear